OpenStudy (anonymous):

Horizontal asymptote of the function f(x)= 6x^3+17x^2-528/2x^3-14x+321

OpenStudy (anonymous):

\[f(x) = 6x^3+17x^2-528/2x^3-14x+321\]

OpenStudy (anonymous):

can someone help me please

OpenStudy (callisto):

Horizontal asymptotes \[=f(x) = \lim_{x \rightarrow \infty} \frac{6x^3+17x^2-528}{2x^3-14x+321}=...\]

OpenStudy (anonymous):

yes thats the problem

OpenStudy (callisto):

Divide both the numerator and denominator by x^3

OpenStudy (callisto):

And \[\lim_{x \rightarrow \infty} \frac{1}{x} =0\]

OpenStudy (anonymous):

i got 3

OpenStudy (callisto):

Isn't it correct?!

OpenStudy (anonymous):

what is the next step , i believe 3 is right

OpenStudy (callisto):

It is 3..

OpenStudy (callisto):

\[\lim_{x \rightarrow \infty} \frac{6x^3+17x^2-528}{2x^3-14x+321}\]\[=\lim_{x \rightarrow \infty} \frac{\frac{6x^3+17x^2-528}{x^3}}{\frac{2x^3-14x+321}{x^3}}\]\[=\lim_{x \rightarrow \infty} \frac{6+\frac{17}{x}-\frac{528}{x^3}}{2-\frac{14}{x^2}+\frac{321}{x^3}}\]Then take the limit